OpenStudy (anonymous):

The volume of a cylinder is 72 ft3. The volume of a cube is 36 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube? Provide an explanation for your answer to receive full credit.

OpenStudy (cwrw238):

The following should help: Volume of a cylinder = pi r^2 h .. .. .. . cone = (1/3) pi r^2 h Volume of a cube = L^3 where L = length of each side Volume of a pyramid fitting inside this = (1/3)* area of base * L